In woodshop class, you must cut a piece of wood within 3/16 inch of the teacher's specification of 5 1/8 inches. Write an absolute value inequality that describes the acceptable lengths for this piece of wood. Solve the inequality.
Answer by stanbon(75887)

woodshop class, you must cut a piece of wood within 3/16 inch of the teacher's specification of 5 1/8 inches. Write an absolute value inequality that describes the acceptable lengths for this piece of wood. Solve the inequality.
---------------------------------
|x - 5 1/8| <= 3/16
|x - (41/8)| <= 3/16
-3/16 <= x - (41/8) <= 3/16
-3/16 + 41/8 <= x <= 3/16 +41/8
-3/16 + 82/16 <= x <= 3/16 + 82/16
79/16 <= x <= 85/16
===========================================================
Cheers,
Stan H.